The length of a rectangle is 3 feet less than twice the width of the rectangle. If the perimeter of the rectangle is 174 feet, find the width and the length
Answer by nerdybill(7384)

Let w = width of rectangle
then because "length of a rectangle is 3 feet less than twice the width" we have
2w-3 = length of rectangle
.
perimeter of any rectangle is:
2(length+width)
.
2(w + 2w-3) = 174
2w + 4w - 6 = 174
6w - 6 = 174
6w = 180
w = 30 feet (width)
.
length:
2w-3 = 2(30)-3 = 60-3 = 57 feet (length)